# Disentangle and Regularize: Sign Language Production with Articulator-Based Disentanglement and Channel-Aware Regularization

### v1

@@ Abstract
In this work, we propose a simple gloss-free, transformer-based sign language production (SLP) framework that directly maps spoken-language text to sign pose sequences. We first train a pose autoencoder that encodes sign poses into a compact latent space using an articulator-based disentanglement strategy, where features corresponding to the face, right hand, left hand, and body are modeled separately to promote structured and interpretable representation learning. Next, a non-autoregressive transformer decoder is trained to predict these latent representations from sentence-level text embeddings. To guide this process, we apply channel-aware regularization by aligning predicted latent distributions with priors extracted from the ground-truth encodings using a KL-divergence loss. The contribution of each channel to the loss is weighted according to its associated articulator region, enabling the model to account for the relative importance of different articulators during training. Our approach does not rely on gloss supervision or pretrained models, and achieves state-of-the-art results on the PHOENIX14T dataset using only a modest training set.
@@ 1 Introduction
Recent advances in SLP have focused on generating skeletal pose sequences instead of synthesizing sign language videos directly. This intermediate representation provides a structured and computationally efficient way to model sign articulation while separating the challenges of motion synthesis and realistic rendering [19]. Despite progress, generating natural and expressive sign sequences remains an open challenge. Many SLP models tend to produce averaged movements, failing to capture the fine-grained articulations required for natural sign language expressions [30], [17]. This results in overly simplified gestures, particularly in hand and finger movements. Additionally, error accumulation in autoregressive decoding, where errors propagate over time, further degrades sequence coherence, leading to unnatural transitions between signs [20] , [21]. To address these challenges, various techniques have been explored, including adversarial training, progressive transformers, and mixture density networks (MDNs) [22]. While these methods improve realism, they struggle with smooth and expressive sign transitions, particularly in fine hand articulations.
@@
We propose a simple transformer-based SLP framework that directly translates spoken language text into continuous sign pose sequences without intermediate gloss annotations. The contribution of our work to the SLP domain can be summarized as follows:
@@ 3.2 Transformer Model
Figure 2: Overview of the proposed non-autoregressive transformer architecture.
@@ 4.1 Datasets
We conduct our experiments on the PHOENIX-2014T benchmark [3], which consists of German Sign Language (DGS) recordings extracted from weather forecast broadcasts. Specifically, we utilize the 3D pose data provided by the CVPR 2025 SLRTP Challenge [31]. In this dataset, each video is aligned with both gloss-level annotations and spoken German translations. The dataset contains 8,247 sentences with 1,085 unique signs, split into 7,096 training, 519 development, and 642 test instances. 3D pose sequences are obtained by extracting keypoints from videos using MediaPipe Holistic [16]. The 2D predictions are uplifted to 3D following the method proposed by Ivashechkin et al. (2022) [13], resulting in 178 keypoints per frame, formatted as $(N\times K\times D)$ tensors, where $N$ is the number of frames, $K=178$ the number of keypoints, and $D=3$ the spatial dimension.
@@ 4.2.2 Training Generator Model
The training objective of our transformer model is divided into two phase to effectively balance the learning of accurate pose encodings and structured latent space distribution. In the first-phase, the model is trained to predict ground truth (GT) pose encodings obtained from the pretrained pose autoencoder. In this phase, we minimize an L1 loss between the predicted encodings for the transformer model and the obtained GT encodings. We will refer to this model as MP1.
@@
$\mathcal{L}_{\text{enc}}=\sum_{R\in\{\text{B},\text{RH},\text{LH},\text{F}\}}w_{R}\sum_{t=1}^{T}\left\|\hat{z}_{t}^{(R)}-z_{t}^{(R)}\right\|_{1}$ (4)
@@
Since hands carry the core lexical content in sign language, they are given the highest weights ($w_{\text{RH}}=7$, $w_{\text{LH}}=5$). The right hand, in particular, is typically the dominant hand in sign languages like DGS, and is primarily responsible for conveying meaning, while the left hand often plays a supportive or symmetrical role. The body provides broader spatial cues but is less semantically dense per frame, so it has a lower weight ($w_{\text{B}}=1$). The face, while critical for grammatical and emotional expressions, involves highly redundant motion across many keypoints, leading to a moderate weight ($w_{\text{F}}=2$). This design ensures the model focuses on the most informative regions for sign language generation.
@@
In the second phase of transformer training, we incorporate a KL divergence term to promote generalization and preserve the distributions of the structurally disentangled latent space. We refer to the resulting model as MP2, which extends the first-phase model (MP1) by adding regularization on the prior channel distributions. In this context, during training, we enforce the predicted latent channel distributions to align with precomputed priors derived from ground truth encodings for each articulator region. This loss term is provided in Eqn. (7).
@@
Here, $\mu^{(c)}$ and $\sigma^{(c)}$ denote the batch-wise mean and variance of the predicted encoding channels, computed dynamically at each training iteration, while $\mu_{prior}^{(c)}$ and $\sigma_{prior}^{(c)}$ represent the prior statistics derived from the training data embeddings. Here, $R$ is the similar region set $\{B,RH,LH,F\}$ that we explained previously, and each channel is weighted according to the region it belongs to. In our experiments we set the weights as $\lambda_{\text{B}}=1$, $\lambda_{\text{RH}}=5$, $\lambda_{\text{LH}}=3$, $\lambda_{\text{F}}=1$.
@@ 5 Discussions
5.1 Analysis of the Learned Latent Space
@@
As can be seen, the face channels have minimal variance, low interquartile ranges, and sharply peaked histograms, indicating high interdependence among the facial keypoints; hence lower average entropy (i.e. 0.1377). This justifies our decision to allocate a smaller latent capacity to the face despite its high input dimensionality, as much of the motion can be captured through a few dominant modes. In contrast, the right and left hand channels show broader, more dispersed distributions with higher entropy and standard deviation across most dimensions, confirming the rich variability and critical role of manual articulators in sign expression. The body encodings fall somewhere in between, reflecting more stable but still semantically relevant movement, however body region spans a larger physical space and thus generates higher magnitude latent activations, even when the underlying motion is less semantically dense. To mitigate this effect and maintain balanced learning, we increase the contribution of semantically richer regions by weighting.
These distributional differences not only validate the semantic partitioning of the latent space but also reinforce the need for targeted regularization and loss weighting, encouraging the model to prioritize dense, information-rich regions while maintaining reconstruction fidelity across all articulators.
5.2 Experimental Results
Table 1: Performance comparison of the models on the PHOENIX14T Test Set.
Models Gloss BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE chrF WER
@@ 5.2 Experimental Results
PT Base, [30] 6.27 3.33 2.14 1.59 14.52 9.50 96.86
PT FP&GN, [30] 11.45 7.08 5.08 4.04 19.09 14.52 94.65
PoseVQ-MP [33] 15.43 10.69 8.26 6.98 - - -
PoseVQ-DDM [33] 16.11 11.37 9.22 7.50 - - -
SignVQNet [12] - - - 6.88 - - -
Stitching T2P (Continuous) [30] 25.14 13.54 8.98 6.67 29.5 26.49 -
GCDM [26] 22.03 14.21 10.16 7.91 23.20 - 81.94
@@
NAT-AT [8] 14.26 9.93 7.11 5.53 18.72 - 88.15
NAT-EA [8] 15.12 10.45 7.99 6.66 19.43 - 82.01
NSVQ+Non-Autoreg. Decoding [29] 27.74 16.36 11.75 9.20 27.93 - -
GT - 34.43 22.08 16.13 12.81 35.22 34.62 85.81
Ours (MP1) 27.14 15.31 10.47 7.98 27.92 27.60 90.18
Ours (MP2) 31.42 18.87 13.17 10.02 30.90 30.67 88.44
Our experimental results, along with comparisons to state-of-the-art methods, are presented in Table 1. As shown in the table, early training alone (i.e., the MP1 model) achieves strong back-translation performance, surpassing several existing models and demonstrating the effectiveness of utilizing ground truth encodings from the autoencoder.
In the second phase, the inclusion of KL divergence regularization proves critical in mitigating overfitting to dominant high-variance patterns, such as higher range body movements, prevalent in the training data. This regularization encourages smoother outputs and better alignment with the underlying distributions of the disentangled latent regions.
This two-phased training strategy directly contributes to improved generalization, as evidenced by the enhanced back-translation scores in Table 1. Notably, applying KL divergence regularization on top of the MP1 model results in a substantial performance gain, particularly on the test set. The effect of this regularization is clearly observed when comparing MP1 and MP2 (i.e., MP1 with added KL divergence). The BLEU-4 score increases from 7.98 (MP1) to 10.02 (MP2), while ROUGE improves from 27.92 to 30.90, highlighting the effectiveness of aligning predicted latent distributions with learned priors during training.
We empirically observed that applying KL divergence regularization directly in the first phase of the training leads to suboptimal performance, constraining the latent space before it could capture useful variation. Instead, we introduced the KL term after an initial reconstruction learning phase, which leads to improved generalization. This behavior parallels the trends observed in VAE-based sequence models, where KL regularization applied on a premature model often causes latent representations to collapse toward the prior, reducing their utility [2]. To address this, prior work has proposed staged or cyclical KL annealing strategies [25, 6]. Although our model does not learn a variational posterior, our findings suggest that latent space regularization benefits from careful scheduling of KL divergence even in deterministic, structured representation settings.
Since our model operates without gloss supervision, the reported Word Error Rate (WER) should be interpreted with caution. Notably, even when evaluating on ground-truth pose sequences, our back-translation model yields a WER of 85.81. In contrast, gloss-based models [8, 26] achieve considerably lower WERs on the same ground-truth data (e.g., 55.93 and 74.17, respectively). This discrepancy underscores a key limitation of using WER as an evaluation metric in gloss-free SLP settings. This highlights the limitation of WER in gloss-free SLP scenarios and emphasizes the need for evaluation metrics that assess generation quality directly within the continuous pose space, rather than depending on intermediate textual representations.

### v2

@@ Abstract
In this work, we propose DARSLP, a simple gloss-free, transformer-based sign language production (SLP) framework that directly maps spoken-language text to sign pose sequences. We first train a pose autoencoder that encodes sign poses into a compact latent space using an articulator-based disentanglement strategy, where features corresponding to the face, right hand, left hand, and body are modeled separately to promote structured and interpretable representation learning. Next, a non-autoregressive transformer decoder is trained to predict these latent representations from sentence-level text embeddings. To guide this process, we apply channel-aware regularization by aligning predicted latent distributions with priors extracted from the ground-truth encodings using a KL-divergence loss. The contribution of each channel to the loss is weighted according to its associated articulator region, enabling the model to account for the relative importance of different articulators during training. Our approach does not rely on gloss supervision or pretrained models, and achieves state-of-the-art results on the PHOENIX14T and CSL-Daily datasets.
@@ 1 Introduction
Recent advances in SLP have focused on generating skeletal pose sequences instead of synthesizing sign language videos directly. This intermediate representation provides a structured and computationally efficient way to model sign articulation while separating the challenges of motion synthesis and realistic rendering [19]. Despite progress, generating natural and expressive sign pose sequences remains an open challenge. Many SLP models tend to produce averaged movements, failing to capture the fine-grained articulations required for natural sign language expressions [30], [17]. This results in overly simplified gestures, particularly in hand and finger movements. Additionally, error accumulation in autoregressive decoding, where errors propagate over time, further degrades sequence coherence, leading to unnatural transitions between signs [20] , [21]. To address these challenges, various techniques have been explored, including adversarial training, progressive transformers, and mixture density networks (MDNs) [22]. While these methods improve realism, they struggle with smooth and expressive sign transitions, particularly in fine hand articulations.
@@
To address these challenges, we propose DARSLP, a simple transformer-based SLP framework that directly translates spoken language text into continuous sign pose sequences without gloss supervision. The contribution of our work to the SLP domain can be summarized as follows:
@@ 3.2 Transformer Model
Figure 2: Overview of the proposed DAR-SLP architecture.
@@ 4.1 Datasets
We evaluate our model on two continuous sign language datasets. PHOENIX-2014T [3] contains 8,247 German Sign Language sentences aligned with gloss and spoken German. We use the 3D pose annotations from the CVPR 2025 SLRTP Challenge [31], where 2D keypoints extracted via MediaPipe Holistic [16] are uplifted to 3D using Ivashechkin et al. ’s method [13], yielding $N\times 178\times 3$ tensors. CSL-Daily [35] includes 20,654 Chinese Sign Language (CSL) sentences from 10 signers. We extract 3D poses using MediaPipe Holistic, standardizing each frame to 178 keypoints (upper body, hands, face). To ensure consistency across signers and datasets, poses are normalized by centering on the neck and scaling by shoulder width.
@@ 4.2.2 Training Generator Model
The training objective of our transformer model is divided into two phase to effectively balance the learning of accurate pose encodings and structured latent space distribution. In the first-phase, the model is trained to predict ground truth (GT) pose encodings obtained from the pretrained pose autoencoder. In this phase, we minimize an L1 loss between the predicted encodings for the transformer model and the obtained GT encodings. We will refer to this model as DARSLP.
@@
$\mathcal{L}_{\text{enc}}=\sum_{R\in\{\text{B},\text{RH},\text{LH},\text{F}\}}w_{R}\frac{1}{T}\sum_{t=1}^{T}\left\|\hat{z}_{t}^{(R)}-z_{t}^{(R)}\right\|_{1}$ (4)
@@
Since hands carry the core lexical content in sign language, they are given the highest weights ($w_{\text{RH}}=7$, $w_{\text{LH}}=5$). The right hand, in particular, is typically the dominant hand in sign languages like DGS and CSL, and is primarily responsible for conveying meaning, while the left hand often plays a supportive or symmetrical role. The body provides broader spatial cues but is less semantically dense per frame, so it has a lower weight ($w_{\text{B}}=1$). The face, while critical for grammatical and emotional expressions, involves highly redundant motion across many keypoints, leading to a moderate weight ($w_{\text{F}}=2$). This design ensures the model focuses on the most informative regions for sign language generation.
@@
In the second phase of transformer training, we incorporate a KL divergence term to promote generalization and preserve the distributions of the structurally disentangled latent space. We refer to the resulting model as DARSLP-KL, which extends the first-phase model by adding regularization on the prior channel distributions. In this context, during training, we enforce the predicted latent channel distributions to align with precomputed priors derived from ground truth encodings for each articulator region. This loss term is provided in Eqn. (7).
@@
Here, $\mu^{(c)}$ and $\sigma^{(c)}$ denote the batch-wise mean and variance of the predicted encoding channels, computed dynamically at each training iteration, while $\mu_{prior}^{(c)}$ and $\sigma_{prior}^{(c)}$ represent the prior statistics derived from the training data embeddings. Here, $R$ is the similar region set $\{B,RH,LH,F\}$ that we explained previously, and each channel is weighted according to the region it belongs to. For initial experiments we set the weights as $\lambda_{\text{B}}=1$, $\lambda_{\text{RH}}=5$, $\lambda_{\text{LH}}=3$, $\lambda_{\text{F}}=1$.
@@
4.3 Ablations
Ablation study on latent dimensionality. To assess the impact of latent dimensionality on sign pose generation, we conduct an ablation study by varying the allocation of dimensions across body, hands, and face components. These experiments test whether increased capacity improves downstream performance and validate our design choices. We compare three configurations: (1) 80 dimensions (64 for body and hands, 16 for face), (2) 96 dimensions (64 + 32), and (3) 160 dimensions (128 + 32). Within each, body and hand dimensions are distributed proportionally to joint counts to maintain anatomical consistency.
As shown in Table 1, the 80-dimensional setup achieves the best overall performance. Increasing face capacity to 32 dimensions shows no gain, while expanding body and hand dimensions to 128 degrades results. The ablation further validates our face encoding decision. We demonstrated that the 128×3 facial keypoints, due to their strong inter-keypoint dependency, can be effectively compressed to just 16 latent dimensions without compromising accuracy. This structural redundancy in facial movements, confirmed by our variance analysis enables robust and efficient encoding (see Section 5.1). Increasing the dimensionality beyond this point introduces unnecessary capacity without a gain, confirming the compact configuration as both sufficient and optimal.
Table 1: Performance comparison of the models with different latent dimensions on the PHOENIX14T DEV Set. All models are trained without KL divergence.
Models DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
GT 0.000 30.92 21.34 16.01 12.74 34.71 30.19
128 (Body+Hands) + 32 (Face) 0.042 22.45 13.78 9.58 7.21 28.18 20.70
64 (Body+Hands) + 32 (Face) 0.044 20.92 12.57 8.79 6.72 27.78 19.91
64 (Body+Hands) + 16 (Face) 0.041 22.59 13.87 9.83 7.59 28.25 21.19
Ablation study on generator training. To determine the optimal setup for our transformer-based generator, we conducted ablation studies on training strategies and region-wise weighting. We initially used moderate region weights (1–7–5–2) for the L1 loss and (1–3–5–1) for the KL divergence, corresponding to the body, right hand, left hand, and face regions, respectively, to reflect articulator importance without imposing a strong bias. This served as a baseline for assessing the impact of KL regularization and its scheduling. As shown in Table 2, omitting KL leads to suboptimal performance (BLEU-4: 7.59), while applying it from the start causes instability (BLEU-4: 6.93). Our two-stage training strategy, introducing KL after initial training, yields the best results (BLEU-4: 8.70). This aligns with findings in VAE-based models, where premature KL application can collapse the latent space toward the prior [2], prompting staged or cyclical KL schedules [25, 6]. Though our model does not learn a variational posterior, our findings suggest that latent space regularization benefits from careful scheduling of KL divergence even in deterministic, structured representation settings.
After establishing the training pipeline with this balanced setup, we performed further ablations focusing on region-specific weighting. These experiments explore how emphasizing specific articulators (particularly the hands) in both L1 and KL terms influences the generation quality (Tables 3 and 4).
Table 2: Performance comparison of the models with different training strategies on the PHOENIX14T DEV Set. All models use 80-dimensional latent space.
Models DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
GT 0.000 30.92 21.34 16.01 12.74 34.71 30.19
L1 w/o KL 0.041 22.59 13.87 9.83 7.59 28.25 21.19
L1 + w/ KL 0.042 20.51 12.46 8.88 6.93 27.00 18.95
Pretrained(w/L1)+Tuning w/ KL 0.039 25.91 16.34 11.43 8.70 29.92 23.67
Table 3: Performance comparison of models with different region weight configurations on the PHOENIX14T DEV set. All models use 80-dimensional latent space, and are trained without KL regularization.
Weighting Strategy (Body–RH–LH–Face) DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
Uniform (1–1–1–1) 0.039 22.14 13.14 8.95 6.72 27.89 19.91
Moderate Hand Emphasis (1–7–5–2) 0.041 22.59 13.87 9.83 7.59 28.25 21.19
Heavy Hand Emphasis (1–14–10–2) 0.043 24.78 15.84 11.26 8.66 29.36 22.98
Table 4: Ablation results with different KL divergence weights with fixed L1 weighting (1–14–10–2) on the PHOENIX14T DEV set. All models use 80-dimensional latent space, and are trained with staged KL regularization.
KL Weights (Body–RH–LH–Face) DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
Uniform (1–1–1–1) 0.040 26.69 17.49 12.79 9.99 30.63 24.92
Moderate Hand Emphasis (1–3–5–1) 0.040 27.35 17.51 12.52 9.66 30.96 25.28
Heavy Hand Emphasis (1–10–6–1) 0.039 27.11 18.00 13.09 10.27 31.38 25.70
Based on the ablation results, region-wise L1 loss weights are set to (1–14–10–2), and KL divergence is introduced in the second phase with weights (1–10–6–1).
@@ 5 Discussions
5.1 Analysis of the Learned Latent Spaces
@@
As can be seen, the face channels have minimal variance. In contrast, the right and left hand channels show broader, more dispersed distributions with higher entropy and standard deviation across most dimensions, confirming the rich variability and critical role of manual articulators in sign expression. The body encodings fall somewhere in between, reflecting more stable but still semantically relevant movement, however body region spans a larger physical space and thus generates higher magnitude latent activations, even when the underlying motion is less semantically dense. To mitigate this effect and maintain balanced learning, we increase the contribution of semantically richer regions by weighting.
These distributional differences confirm that the latent space is semantically partitioned. They also highlight the need for targeted regularization and proper loss weighting. This encourages the model to focus more on dense, information-rich regions while still preserving reconstruction quality across all articulators.
(a) Right hand channels (avg. entropy: 2.98).
(b) Left hand channels (avg. entropy: 3.19).
(c) Face channels (avg. entropy: 2.21).
(d) Body channels (avg. entropy: 1.65).
Figure 4: Histograms of top 8 latent channels with the highest entropy for each region of CSL-Daily Dataset. Each subplot shows the distribution of a channel along with its entropy (E), interquartile range (IQR), and standard deviation (SD). (Zoom in for better visibility.)
The latent space structure observed in CSL-Daily mirrors that of PHOENIX14T, with clear separation between articulator groups and similar relative entropy patterns, manual articulators exhibit the highest variability, followed by face and body (Figure 4). CSL-Daily encodings exhibit broader distributions across all articulator groups, as reflected by higher entropy values, particularly in the manual articulators. The left and right hand channels display the widest spread, with average entropies around 3.0, indicating increased variability and expressiveness in CSL’s daily conversational context. Even face channels, which previously showed limited variance, now present more active and information-rich distributions. This broader activation aligns with CSL-Daily’s more diverse signer pool and open-domain language, reinforcing the need for careful latent space design.
5.2 Test Results
Table 5: Performance comparison of the models on the PHOENIX14T TEST Set. †  denotes that the scores are obtained using the back translation model shared in SLRTP-Sign-Production-Evaluation [31].
Models Gloss DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ ROUGE↑ chrF↑
GT - 0.000 34.43 22.08 16.13 12.81 35.22 34.62
@@ 5.2 Test Results
PT Base [30] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
PT FP&GN [30] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
G2P-MP [33] 0.146 15.43 10.69 8.26 6.98 - -
G2P-DDM [33] 0.116 16.11 11.37 9.22 7.50 - -
SignVQNet [12] 0.299 - - - 6.88 -
Stitching T2P (Continuous) [30] 0.572 25.14 13.54 8.98 6.67 29.5 26.49
GCDM [26] - 22.03 14.21 10.16 7.91 23.20 -
@@
NSLP-G [12] 0.638 - - - 5.56 - -
NSLP-G w/o Finetuning [12] 0.646 - - - 4.41 - -
NAT-AT [8] 0.177 14.26 9.93 7.11 5.53 18.72 -
NAT-EA [8] 0.146 15.12 10.45 7.99 6.66 19.43
NSVQ+Non-Autoreg. Decoding [29] 0.105 27.74 16.36 11.75 9.20 27.93 -
Ours-DARSLP 0.043 28.91 16.45 11.52 8.86 28.50 28 .84
Ours-DARSLP-KL 0.039 32.33 19.84 14.07 10.90 31.80 32.00
We evaluate the best model configuration on the PHOENIX14T and CSL-Daily test sets. Table 5 presents PHOENIX14T results alongside prior work. The DARSLP model, already outperforms several baselines, demonstrating the effectiveness of using ground-truth encodings from the autoencoder. Adding KL divergence in the second phase significantly improves generalization by reducing overfitting to dominant high-variance patterns (e.g., body motion) and encouraging smoother outputs aligned with latent region distributions. This yields measurable gains: BLEU-4 improves from 8.86 to 10.90 and ROUGE from 28.50 to 31.80, confirming the benefit of aligning predicted embeddings with learned priors.
@@
Table 6: Evaluation results on the CSL-Daily TEST Set using the finalized configuration.
Model DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ ROUGE↑ chrF↑
GT 0.000 21.63 10.90 6.31 4.02 21.81 5.26
DARSLP 0.177 19.13 8.33 4.46 2.73 18.72 4.24
DARSLP-KL 0.176 20.03 9.16 5.25 3.35 19.41 4.61
Generalization to CSL-Daily. We evaluate our finalized transformer model, selected on PHOENIX14T, on the CSL-Daily dataset using the same architecture and training strategy. To accommodate CSL-Daily’s broader linguistic context and higher variance, we added another projection layer when mapping keypoints to the 80-dimensional latent space. This produced latent distributions, particularly for the right hand, that aligned more closely with the assumed Gaussian prior, making KL regularization more effective (see Supplementary Material for details). As shown in Table 6, the model with KL divergence regularization outperforms the variant trained without KL across all metrics. Despite the domain shift, our model achieves competitive BLEU and ROUGE scores relative to ground-truth, which sets an upper-bound, demonstrating strong generalization to new language content and signer variability. As no prior SLP results exist for sign pose generation task on this dataset, we compare our performance against the ground truth (GT) baseline.